OpenStudy (anonymous):

Tony is drawing a regular polygon. He has drawn two of the sides with an interior angle of 120°, as shown below. What will be the sum of the interior angles of the regular polygon have when Tony completes his drawing?

OpenStudy (anonymous):

each interior angle =180 -360/n (n number of sides) here 120=180 -360/n or 360/n =60 hence n=6 the sum of the interior angles of the regular polygon have when Tony completes his drawing =180((6-2) =720